One number is 20 more than the other. If their sum is 35, what are the numbers?
Answer by math_helper(2461)


7.5 and 27.5
You should be able to do this in your head.

Algebraic work-out:
x + (x+20) = 35
2x + 20 = 35
2x = 15
x = 7.5, x+20 = 27.5